OpenStudy (anonymous):

3. The area of a rectangular swimming pool is 10x^2 – 19x – 15. The length of the pool is 5x + 3. What is the width of the pool?

OpenStudy (anonymous):

2x – 18 2x – 5 5x – 5 5x – 22

OpenStudy (phi):

Area = L*W so W= Area/L (divide both sides by L) Here area is 10x^2 – 19x – 15 divide it by L=5x + 3 Do you know how to divide 5x+3 into 10x^2 – 19x – 15 to get W

OpenStudy (campbell_st):

it factorises (5x +3)(2x -5) if the length is 5x + 3 the width is 2x - 5

OpenStudy (anonymous):

thanks